At an atmospheric pressure of 1.01×105Pa, carbon dioxide (CO2) sublimes at a temperature of -78.5∘C. The heat of sublimation is 5.71×105J/kg. You let heat flow into 6.0 kg of solid CO2 (volume 4.0×10−3m3) at atmospheric pressure and a constant -78.5∘C until it has completely sublimed to gas at -78.5∘C.The CO2 gas occupies a volume of 3.4 m3.

A) For this process, calculate how much heat flows into the CO2.

B) For this process, calculate how much work the CO2 does

C) For this process, calculate the internal energy change of the CO2
Answer #1

Solution:

A) Given,

the heat of sublimation of carbon dioxide h_{s}=5.71\times 10^5 J/Kg .

mass of CO2 (m)= 6.0 Kg

Amount of heat flows into CO2 is given by;

Q= m xh

= 6 x 5.71 x 105 = 3.426 X 106J

B)

Volume of solid CO2 = 4.0 × 10-3 m3

Volume of CO2 gas = 3.4 m3

Therefore, the change in volume on conversion of CO2 is;

AV = 3.4 - (4.0 x 10-3) = 3.396m3

Work done by the CO2 gas against atmospheric pressure is given by;

W=P\Delta V

=1.01\times 10^{5}\times 3.396=3.43\times 10^{5}J

C)

According to the first law of thermodynamics, the change in internal energy is given by

AU = Q-W

  =(3.426\times 10^6)-(3.43\times 10^5)=3.083\times 10^6J